# Adjust for Trust: Mitigating Trust-Induced Inappropriate Reliance on AI Assistance

### v1

@@ 3.2 Experiments
For the ARC task, we recruit users with at least an undergraduate degree. We recruit two groups of 30 users, one each for the ArcC and ArcO settings. Users achieve $67\%$ accuracy on this task without AI assistance, whereas the calibrated and overconfident AI achieve $71\%$ and $64\%$ accuracy, respectively. Users are paid \$1.0, plus a \$0.10 bonus for every correct final decision.
@@
For the Diagnosis task (DiagC task setting), we conduct studies with professional doctors, who achieve $74\%$ task accuracy without the AI. Due to the lower number of qualified participants on Prolific, we recruit 20 users. Users are paid \$2.0, plus a \$0.10 bonus for every correct final decision.
@@
For each task setting, we sample 10 sequences $S_{i}=\{P^{i}_{1},P^{i}_{2},...,P^{i}_{30}\}$ of 30 decision-making problems $P^{i}_{j}=\{(x_{j},y^{*}_{j}),R_{j}\}$. Users in each task setting are randomly assigned to a sequence $S_{i}$ upon starting the study. Appendix A contains additional details about the user study setup.
@@
Figure 4: Reliance metrics at different levels of user trust. In each plot, $r$ represents the weighted Pearson correlation coefficient. All correlations are statistically significant, with $p<0.001$. Bar shades correspond to number of user interactions at each trust level.
Findings.
Figure 4 shows the relationship between user trust and reliance, aggregated across all users in the same task setting. We highlight a few key takeaways (T). T1: Switch Rate is strongly correlated with user trust, which suggests that users’ internal trust influences how likely they are to accept AI advice. T2: User trust has moderate to strong correlation with Over-Reliance. We further observe that Over-Reliance is highest at high values of user trust (999910101010). T3: User trust has a strong negative correlation with Under-Reliance. At lower values of trust ($<5$), users exhibit the most Under-Reliance. These findings suggest that extreme values of user trust act as a cognitive bias, resulting in higher inappropriate reliance in subsequent human-AI interactions.
4 Trust-Adaptive AI Interventions Mitigate Inappropriate Reliance
We hypothesize that AI systems can counterbalance the cognitive bias caused by trust by adapting their behavior according to the user’s trust level. [^3] Our setup assumes that the AI can observe the user’s last-reported trust level at the start of every interaction. We introduce trust-adaptive interventions for reducing inappropriate reliance. Trust-adaptive interventions are designed to correct for trust-induced cognitive bias, and are only applied when the user’s trust level is either above or below a certain threshold. We hypothesize that uniformly applying these interventions, rather than only when trust is low or high, will worsen inappropriate reliance.
Experiment Setup.
We evaluate the utility of adaptive interventions through a between-subjects study. [^4] We do not conduct a within-subjects study because the same user cannot be subjected to multiple conditions without introducing interaction effects. Each user is assigned to one of three experimental conditions: a No Intervention baseline where the intervention is never applied, an Intervention Always baseline where the intervention is applied whenever the AI’s prediction differs from the user’s initial decision, and the Trust-Adaptive Intervention condition where the intervention is only applied when the user’s trust lies above or below a specified threshold. Based on the relationship we observe between reported user trus and reliance (Figure 4), we select a threshold trust of 5 out of 10 for mitigating under-reliance, and 8 out of 10 for mitigating over-reliance. We conduct studies with 30 users assigned to each experimental condition in the ArcC and ArcO task settings, and 20 users for each condition in the DiagC task setting.
Evaluating Interventions.
We evaluate intervention conditions on Under-Reliance or Over-Reliance, depending on the type of inappropriate reliance the intervention is intended to mitigate. We also evaluate Total Inappropriate Reliance, which is the sum of Under-Reliance and Over-Reliance, to check if mitigating one type of inappropriate reliance exacerbates the other type to the same degree. Finally, Final Decision Accuracy captures the effect of interventions on decision-making performance.
We compute metrics across all user interactions where the user and AI disagree ($y^{u,\text{init}}_{i}\neq y^{AI}_{i}$) and also analyze subsets of these interactions based on user trust levels. Aggregating over interactions rather than users may skew user representation, for example, a user with generally low trust will represent more low trust interactions than other users. On the other hand, macro-aggregation by averaging per-user metrics results in high inter-user variance, since some users have very few interactions meeting the specified criteria. Macro-aggregation results are presented in Appendix C.
Figure 5: Reliance metrics and decision accuracy for users, evaluating the utility of supporting explanations at mitigating under-reliance. $n$ represents the number of user-AI interactions that we aggregate over for the corresponding condition. Showing explanations adaptively reduces Under-Reliance and Total Inappropriate Reliance while boosting Final Decision Accuracy across all task settings, particularly when user trust is low.
Figure 6: Reliance metrics and decision accuracy for users, evaluating the utility of counter-explanations at mitigating over-reliance. Showing counter-explanations adaptively reduces Over-Reliance and Total Inappropriate Reliance while boosting Final Decision Accuracy across almost all settings, particularly when trust is high.
4.1 Mitigating Under-reliance with Supporting Explanations
@@
AI explanations have been widely studied as a decision aid Bussone et al. (2015); Wang and Yin (2021); Poursabzi-Sangdeh et al. (2021). Prior work has shown that natural language explanations supporting the AI prediction cause over-reliance Si et al. (2024); Sieker et al. (2024); Hashemi Chaleshtori et al. (2024). We hypothesize that providing natural language supporting explanations when user trust is low ($<5$) can mitigate under-reliance.
@@
We generate supporting explanations for all problems by prompting GPT-4o Hurst et al. (2024) to provide a 3–4 sentence explanation $E^{s}_{i}$ for each option $y_{i}\in\mathcal{Y}$ of each problem $P$ (prompts in Table 5). Explanations were manually reviewed by an author to ensure they entailed the corresponding prediction. When the intervention was applied during a user-AI interaction, the AI would provide an augmented recommendation $R^{\prime}_{i}=R_{i}+E^{s}_{i}$. To encourage users to read the explanation, they are allowed to make their final decision only 15 seconds after the AI advice is shown.
@@
We separately evaluated the interventions across interactions where the user’s initial decision disagrees with the AI prediction, and across the “low trust” subset of interactions. Figure 6 shows that providing supporting explanations adaptively mitigates under-reliance when user trust is low, especially when AI confidence is miscalibrated.
Trust-Adaptive Explanations Help.
@@
In the Trust-Adaptive condition, users exhibit lower Under-Reliance, lower inappropriate reliance, and higher decision accuracy across all task settings. These improvements are particularly notable when user trust is low and an explanation is provided in the Trust-Adaptive Intervention condition but not in the No Intervention condition. Providing explanations when trust is low results in 13–31% reduction in Under-Reliance, 9–38% reduction in Total Inappropriate Reliance, and 10–19% improvement in Final Decision Accuracy.
@@
The largest improvements occur when the AI system is over-confident, i.e. the ArcO task setting, which had the largest number of interactions where user trust was low. This finding indicates that explanations are effective for mitigating AI disuse when AI confidences are miscalibrated.
Persistent Explanations Can Hurt.
The Intervention Always condition does not yield similar improvements and worsens decision accuracy in the Diagnosis task. In the ArcC task setting, showing explanations uniformly increases Under-Reliance when trust is low. We suspect that showing explanations always rather than adaptively exposes the user to many misleading explanations, resulting in an overall loss of trust in the explanations’ trustworthiness. In ArcO and DiagC, explanations reduce inappropriate reliance when trust is low, but not across all user interactions.
Figure 7: Effect of providing supporting explanations (/) and counter-explanations (\), depending on user trust, within the same user session yields complementary benefits in inappropriate reliance and decision-making accuracy.
4.2 Mitigating Over-reliance with Counter-Explanations
@@
Similar to how providing supporting explanations are an effective intervention when user trust is low, we investigate whether providing reasons for why the model prediction might be incorrect can counter-balance the cognitive effect of high trust ($>8$). Such counter-explanations have been shown to reduce over-reliance compared to regular supporting explanations Si et al. (2024).
@@
Similar to the supporting explanations, we generate natural language counter-explanations for each option by prompting GPT-4o to list 1–2 reasons why that option might be incorrect, while not completely rejecting that option (e.g. “I believe Bronchitis is the correct diagnosis due to …, but it is possible that…”). The counter-explanations frequently include expressions of uncertainty (“could potentially”, “it may be that”), alternative possibilities, and specific circumstances under which the model’s prediction may be incorrect. Table 6 contains examples of generated counter-explanations.

### v2

@@ 4.3. Participant Recruitment
For the ARC task, we recruit participants on the Prolific platform who have at least an undergraduate degree. Users achieve $67\%$ accuracy on this task without AI assistance. Users are paid a base amount of \$1.0, with a bonus of \$0.10 per correct final decision. For each ARC task setting in our experiments, we recruit 30 users per experimental condition.
@@
For the Diagnosis task, we conduct studies with professional medical doctors recruited through Prolific. These users achieve $74\%$ task accuracy without the AI advice. [^2] We reject study data from users whose initial decision accuracy is near random chance ($25\%$), since they may have falsified their qualifications on Prolific. Users are paid \$2.0 (since the Diagnosis task takes longer), plus a \$0.10 bonus for every correct final decision. Due to the lower number of qualified participants on Prolific, we recruit 20 users per condition in our Diagnosis experiments.
4.4. Task Settings and Sampling Sequences of Decision-Making Problems.
We perform user studies with three task settings: the ARC task with a calibrated AI (ArcC), the ARC task with an overconfident AI (ArcO), and the Diagnosis task with a calibrated AI (DiagC). [^3] We do not conduct experiments on the Diagnosis task with an overconfident AI due to being unable to recruit sufficient participants on Prolific. For each task setting, we sample 10 sequences $S_{i}=\{P^{i}_{1},P^{i}_{2},...,P^{i}_{30}\}$ of 30 decision-making problems $P^{i}_{j}=\{(x_{j},y^{*}_{j}),R_{j}\}$, where each problem is also accompanied by an AI recommendation. In these sampled problem sequences, the calibrated and overconfident AI assistants have an accuracy of $71\%$ and $64\%$, respectively. Figure 4 shows calibration curves for the calibrated and overconfident AI.
@@
Users in each task setting are randomly assigned to a sequence $S_{i}$ upon starting the study. Appendix B contains additional details about the user study setup.
Figure 4. Calibration curves and Expected Calibration Error (ECE) of our simulated AI assistants.
Calibration curves and Expected Calibration Error (ECE) of our simulated AI assistants.
4.5. Measuring Users’ Reliance Behavior
We inspect interactions where the user’s initial decision differs from the AI prediction, i.e. $y^{u,\text{init}}_{i}\neq y^{AI}_{i}$. We capture the degree of users’ reliance on AI assistance using Switch Rate (Yin et al., 2019; Zhang et al., 2020), the fraction of interactions where the user switched their decision to the AI prediction. Following Ma et al. (2024); Swaroop et al. (2025), we capture the degree of appropriate reliance on AI assistance using two metrics: Over-Reliance represents the fraction of AI-incorrect interactions where the user switches to the AI’s prediction, while Under-Reliance represents the fraction of AI-correct interactions where the user does not switch to the AI’s prediction.
5. Preliminary Study: How does User Trust Impact Reliance on AI Assistance?
In our first experiment, we investigate the correlation between users’ self-reported trust levels [^4] Henceforth, “user trust” refers to the user’s trust level at the start of an interaction, i.e. the trust score they reported at the end of the previous interaction. and their reliance behavior in subsequent interactions. Previous works have investigated whether users’ personality traits, specifically their propensity to trust, are correlated with their reliance on AI assistance (Riedl, 2022; Swaroop et al., 2024); these works have found that these personality traits are not predictive of reliance behavior. Swaroop et al. (2025) instead try to determine a user’s likelihood of over-relying in future interactions by using probe questions. In contrast, we hypothesize that the trust level reported by the user at the end of the $i^{th}$ decision-making problem is correlated with their likelihood of accepting AI recommendations during the $(i+1)^{th}$ user-AI interaction.
  1. H1:
    When the user’s initial decision disagrees with the AI prediction, their likelihood of accepting the AI prediction (captured by the Switch Rate) is correlated with the trust level they reported at the end of the previous decision-making problem.
We further hypothesize that when user trust is either low or high, users make worse decisions about whether to rely on AI recommendations or not in the next user-AI interaction. When user trust in the AI assistant is low (i.e., they do not believe the AI assistant can help them with the decision-making task), they are more likely to disregard correct AI recommendations. Similarly, when user trust is high, they are more likely to accept incorrect AI recommendations.
  1. H2:
    Users will have higher Over-Reliance when their trust in the AI assistant at the start of the interaction is high.
  2. H3:
    Users will have higher Under-Reliance when their trust in the AI assistant at the start of the interaction is low.
5.1. Experiment Setup
We conducted user studies across all three task settings (ArcC, ArcO and DiagC), recruiting 30 participants for each of the ARC task settings and 20 participants for the DiagCsetting. For each task setting, we aggregate all user-AI interactions across all participants, excluding the very first interaction for each participant since we do not have a user-reported trust level. We further exclude all interactions where the user’s initial decision $y^{u,\text{fin}}_{i}$ matches the AI prediction $y^{AI}_{i}$, since in these interactions users almost always stick with their original decision. Finally, we bin the remaining interactions in each task setting by the user’s last reported trust level, and compute Switch Rate, Under-Reliance and Over-Reliance for each trust bin. We compute a weighted Pearson correlation coefficient between user trust levels and the above reliance metrics, where trust bins are weighted by the number of corresponding user-AI interactions.
@@ 5.1. Experiment Setup
Figure 5. Reliance metrics at different levels of user trust. In each plot, $r$ represents the weighted Pearson correlation coefficient. All correlations are statistically significant, with $p<0.001$. Bar shades correspond to number of user interactions at each trust level.
Reliance metrics at different levels of user trust.
5.2. Results
Figure 5 shows the relationship between user trust and reliance, aggregated across all users in the same task setting. We highlight a few key takeaways:
Switch Rate is strongly correlated with user trust (hypothesis H1):
Correlations between user trust and Switch Rate vary from 0.7960.7960.9490.949, indicating a strong correlation between users’ trust levels and their likelihood of accepting AI assistance. These correlations suggest that users’ internal trust influences how likely they are to accept AI advice.
Users have higher Over-Reliance when their trust level is high (hypothesis H2):
User trust has a moderate to strong correlation (0.5000.5000.7800.780) with Over-Reliance. We further observe that Over-Reliance is highest at high values of user trust (991010).
Users have higher Under-Reliance when their trust is low (hypothesis H3):
User trust has a strong negative correlation (0.903-0.903 to 0.742-0.742) with Under-Reliance. At lower values of trust (<5<5), users exhibit highest Under-Reliance.
Taken together, these findings suggest that extreme values of user trust act as a cognitive bias for users in subsequent human-AI interactions, resulting in higher inappropriate reliance.
5.3. Motivating Trust-Adaptive AI Interventions for Mitigating Inappropriate Reliance
Our findings verify our hypothesis that users’ self-reported trust levels (representing the user’s belief in the AI’s helpfulness at the start of the interaction) is correlated with their reliance behavior in the next interaction. Further, they suggest that when user trust is at extreme levels (either low or high), it acts as a cognitive bias that clouds user judgment when deciding whether to rely on future AI recommendations or not, resulting in inappropriate reliance. In a scenario where AI assistants can observe a user’s trust level after each interaction, we investigate whether the AI can counterbalance this trust-induced cognitive bias by adapting its behavior the next time it provides a recommendation to that user. For instance, when the user has low trust and is more likely to disregard AI assistance, the assistant can encourage the user to consider the recommendation more closely by providing extra information to support its prediction. We introduce trust-adaptive interventions for reducing inappropriate reliance. Trust-adaptive interventions are designed to correct for trust-induced cognitive bias, and are only applied when the user’s trust level at the start of an interaction is either above or below a certain threshold. Our experiments in the following sections will demonstrate the utility of variance trust-adaptive interventions for mitigating trust-induced inappropriate reliance.
6. Experiment #1: Mitigating Under-reliance with Supporting Explanations
@@
AI explanations have been widely studied as a decision aid (Bussone et al., 2015; Wang and Yin, 2021; Poursabzi-Sangdeh et al., 2021). Prior work has shown that natural language explanations supporting the AI prediction cause over-reliance (Si et al., 2024; Sieker et al., 2024; Hashemi Chaleshtori et al., 2024). Specifically, it has been shown that when users are provided long-form text explanations, they are prone to believing plausible-sounding explanations without closely examining their accuracy, logical consistency, and reasoning coherence. However, when user trust is low, we hypothesize that providing explanations to users will mitigate undesirable disregarding of AI recommendations, while also enabling more critical examination of model explanations.
  1. H4:
    Providing natural language supporting explanations when user trust is low will lead to lower Under-Reliance without an equal increase in Over-Reliance, resulting in an overall increase in users’ final decision accuracy compared to not adapting AI behavior in response to user trust.
  2. H5:
    Providing support explanations adaptively during moments of low trust will yield lower Under-Reliance and higher decision accuracy than providing them during all interactions, regardless of users’ trust levels.
6.1. Experiment Setup
6.1.1. Conditions
We evaluate the utility of adaptive interventions through a between-subjects study. [^5] We do not conduct a within-subjects study because the same user cannot be subjected to multiple conditions without introducing interaction effects. Each user is assigned to one of three experimental conditions:
  1. (1)
    No Intervention: The AI’s behavior stays the same (i.e. it always provides a prediction and a numeric confidence score) throughout the user study, regardless of the user’s trust level.
  2. (2)
    Supporting Explanation Always Shown: The supporting explanation is provided to the user alongside the AI prediction and confidence, regardless of the user’s trust level.
  3. (3)
    Trust-Adaptive Supporting Explanations: The supporting explanation is only provided when the user’s trust lies below a specified threshold. Based on the relationship we observe between reported user trust and reliance (Figure 5), we select a threshold of 5 out of 10, such that explanations are shown whenever the user’s trust level is at 4 or lower.
Similar to our experiments in Section 5, we conduct studies with 30 users assigned to each experimental condition in the ArcC and ArcO task settings, and 20 users for each condition in the DiagC task setting.
6.1.2. Generating and Providing Supporting Explanations
@@
We generate supporting explanations for all problems by prompting GPT-4o (Hurst et al., 2024) to provide a 3–4 sentence explanation $E^{s}_{i}$ for each option $y_{i}\in\mathcal{Y}$ of each problem $P$ (prompts in Table 6). Explanations were manually reviewed by an author to ensure they entailed the corresponding prediction. When the intervention was applied during a user-AI interaction, the AI would provide an augmented recommendation $R^{\prime}_{i}=R_{i}+E^{s}_{i}$. To encourage users to read the explanation, they are only allowed to make their final decision 15 seconds after the AI advice is shown.
6.1.3. Evaluating Interventions.
We evaluate the different experimental conditions on which yields the lowest Under-Reliance by users. To evaluate whether mitigating under-reliance simultaneously exacerbates over-reliance to the same degree, we also evaluate Total Inappropriate Reliance, which is the sum of Under-Reliance and Over-Reliance. Finally, Final Decision Accuracy captures the effect of interventions on users’ decision-making performance.
Similar to Section 5, we compute metrics across all user interactions where the user and AI disagree ($y^{u,\text{init}}_{i}\neq y^{AI}_{i}$) and analyze subsets of these interactions based on user trust levels. Aggregating over interactions rather than users may skew user representation; a user with generally low trust will represent more low trust interactions than other users. On the other hand, macro-aggregation by averaging per-user metrics (Appendix C) results in high inter-user variance, since some users have very few interactions meeting the specified criteria.
We separately evaluated the interventions across interactions where the user’s initial decision disagrees with the AI prediction, and across the “low trust” subset of interactions. For the “low trust” interactions, we conduct a clustered bootstrap test (Efron and Tibshirani, 1994) to evaluate significance of our Trust-Adaptive intervention compared to the No Intervention setting. [^6] We do not conduct a bootstrap test when considering all interactions, because for interactions where user trust is not low, AI behavior is identical in the No Intervention and Trust-Adaptive conditions.
Figure 6. Reliance metrics and decision accuracy for users, evaluating the utility of supporting explanations at mitigating under-reliance. $n$ represents the number of user-AI interactions that we aggregate over for the corresponding condition. Showing explanations adaptively reduces Under-Reliance and Total Inappropriate Reliance while boosting Final Decision Accuracy across all task settings, particularly when user trust is low. Asterisks denote significant improvements using a clustered bootstrap test $({{}^{**}:p<0.01;^{***}:p<0.001)}$.
Reliance metrics and decision accuracy for users, evaluating the utility of supporting explanations at mitigating under-reliance.
6.2. Results
@@
Figure 6 shows that providing supporting explanations adaptively mitigates under-reliance when user trust is low, especially when AI confidence is miscalibrated.
Trust-adaptive explanations reduce under-reliance (hypothesis H4).
@@
In the Trust-Adaptive condition, we observe that users exhibit lower Under-Reliance, lower inappropriate reliance, and higher decision accuracy across all task settings. When considering interactions where user trust is low, a clustered bootstrap test (Efron and Tibshirani, 1994) reveals that the majority of improvements of the Trust-Adaptive intervention over the No Intervention condition are statistically significant. Providing explanations when trust is low results in 13–31% reduction in Under-Reliance, 9–38% reduction in Total Inappropriate Reliance, and 10–19% improvement in Final Decision Accuracy.
Persistent explanations can hurt, and are not as useful as trust-adaptive explanations (hypothesis H5).
The Supporting Explanation Always condition does not yield similar improvements and worsens decision accuracy in the Diagnosis task. In the ArcC task setting, showing explanations uniformly increases Under-Reliance when trust is low. We suspect that showing explanations always rather than adaptively exposes the user to many misleading explanations, resulting in an overall loss of trust in the explanations’ trustworthiness. In ArcO and DiagC, explanations reduce inappropriate reliance when trust is low, but not across all user interactions, indicating over-reliance on explanations when trust is not low.
@@
The largest improvements occur when the AI system is over-confident, i.e. in the ArcO task setting, which had the largest number of interactions where user trust was low. This finding indicates that explanations are effective for mitigating miscalibrated AI disuse.
7. Experiment #2: Mitigating Over-reliance with Counter-Explanations
@@
Similar to how providing supporting explanations is an effective intervention when user trust is low, we investigate whether providing reasons for why the AI prediction might be incorrect can counter-balance users’ tendency to readily believe AI outputs when their trust is high. We hypothesize that alerting users to reasons the AI might potentially incorrect may act as a cognitive forcing function (Buçinca et al., 2021), preventing the user from blindly believing AI outputs without due consideration. Such counter-explanations, which present evidence for alternatives apart from the AI prediction, have been shown to reduce over-reliance compared to regular supporting explanations (Cabitza et al., 2023; Si et al., 2024; Cabitza et al., 2024).
  1. H6:
    Adapting AI behavior when user trust is high by providing counter-explanations to the user will reduce Over-Reliance without an equal increase in Under-Reliance, resulting in an overall increase in users’ final decision accuracy to not adapting AI behavior in response to user trust.
  2. H7:
    Providing counter-explanations adaptively during moments of low trust will yield lower Over-Reliance and higher decision accuracy than providing them during all interactions, regardless of users’ trust levels.
7.1. Experiment Setup
7.1.1. Conditions
Similar to §6, each user is assigned to one of three experimental conditions:
  1. (1)
    No Intervention: The AI’s behavior stays the same (i.e., it always provides a prediction and a numeric confidence score) throughout the user study, regardless of the user’s trust level.
  2. (2)
    Counter-Explanation Always Shown: The counter-explanation is provided to the user alongside the AI prediction and confidence score, regardless of the user’s trust level.
  3. (3)
    Trust-Adaptive Counter-Explanations: The supporting explanation is only provided when the user’s trust falls above a specified threshold. We select a threshold trust of 8 out of 10 for applying the trust-adaptive intervention based on the relationship we observe between reported user trust and over-reliance in Figure 5.
7.1.2. Generating and Providing Counter-Explanations
@@
Similar to the supporting explanations, we generate natural language counter-explanations (Table 7) for each option by prompting GPT-4o to list 1–2 reasons why that option might be incorrect, while not completely rejecting that option (e.g., “I believe Bronchitis is the correct diagnosis due to…, but it is possible that…”). The counter-explanations frequently include expressions of uncertainty (“could potentially”, “it may be that”), alternative possibilities, and specific circumstances under which the model’s prediction may be incorrect.
7.1.3. Evaluating Interventions.
Similar to the experiments in §6, we evaluate each experimental condition on users’ Final Decision Accuracy  while also examining whether the two intervention conditions reduces Over-Reliance without increasing the Total Inappropriate Reliance.
Figure 7. Reliance metrics and decision accuracy for users, evaluating the utility of counter-explanations at mitigating over-reliance. Showing counter-explanations adaptively reduces Over-Reliance and Total Inappropriate Reliance while boosting Final Decision Accuracy across almost all settings, particularly when trust is high. Asterisks denote significant improvements using a clustered bootstrap test $(^{*}:p<0.05)$.
Reliance metrics and decision accuracy for users, evaluating the utility of counter-explanations at mitigating over-reliance.
7.2. Results